# Tree-based formulation for the multi-commodity flow problem

### v1

@@ 3.1 Path generation
Thus, finding a column of negative reduced cost for commodity $k$ amounts to computing an $s_{k}$$t_{k}$ path that minimizes $\sum_{e\in p}(c_{e}-\mu_{e})$ and then subtracting the demand dual $\pi_{k}$. Since all edge weights $c_{e}\geq 0$ and $\mu_{e}\leq 0$, the modified costs $c_{e}-\mu_{e}$ remain bounded below, so the pricing problem reduces to a shortest path problem in a weighted graph.
@@ A.2 Runtime details
Instance Tree Path MOSEK Objective
@@
grid1 0.0 0.0 0.0 0.0 - - 8.2732e+05
grid2 0.0 0.1 0.0 0.0 - - 1.7054e+06
grid3 0.0 0.0 0.0 0.0 - - 1.5246e+06
grid4 0.0 0.1 0.0 0.0 - - 3.0317e+06
grid5 0.0 0.1 0.1 0.1 - - 5.0497e+06
grid6 0.3 0.1 0.2 0.1 - - 1.0402e+07
grid7 0.4 0.1 0.3 0.1 - - 2.5864e+07
grid8 1.6 0.2 1.3 0.2 - - 4.1711e+07
grid9 3.4 0.2 2.8 0.2 - - 8.2653e+07
grid10 4.3 0.3 4.1 0.3 - - 1.6411e+08
grid11 3.9 0.2 4.4 0.3 - - 3.2926e+08
grid12 4.4 0.2 4.2 0.3 - - 5.7719e+08
grid13 10.2 0.3 12.4 0.4 - - 1.1593e+09
grid14 5.5 0.3 8.2 0.4 - - 1.8027e+09
grid15 10.6 0.4 18.8 0.6 - - 3.5935e+09
@@
planar30 0.0 0.0 0.0 0.0 - - 4.4351e+07
planar80 0.4 0.1 0.3 0.1 - - 1.8244e+08
planar100 0.3 0.1 0.5 0.1 - - 2.3134e+08
planar150 4.4 0.2 4.1 0.3 - - 5.4809e+08
planar300 1.6 0.2 1.9 0.2 - - 6.8998e+08
planar500 0.7 0.2 1.0 0.2 - - 4.8198e+08
planar800 5.5 0.3 8.6 0.4 - - 1.1674e+09
planar1000 75.3 0.6 53.5 0.9 - - 3.4496e+09
planar2500 5894.7 10.0 866.9 19.3 - - 1.2662e+10
@@
Austin 670.6 14.5 4967.0 48.1 - - 5.4329e+06
Barcelona 0.1 0.1 2.5 0.2 - - 8.8588e+01
BerlinCenter 10.3 0.5 19.0 0.7 - - 2.7529e+07
Birmingham 196.2 3.3 399.3 16.7 - - 2.1565e+05
ChicagoRegional 68.2 6.1 981.8 33.4 - - 5.5131e+06
ChicagoSketch 0.5 0.2 3.1 0.6 - - 6.7059e+06
Philadelphia 380.6 7.4 1471.2 17.3 - - 2.5578e+07
Sydney 66.4 11.1 2520.9 59.7 - - 4.9653e+06
Winnipeg 0.0 0.1 2.9 0.2 - - 2.3767e+02

### v2

@@ 3.1 Path generation
Thus, finding a column of negative reduced cost for commodity $k$ amounts to computing an $s_{k}$$t_{k}$ path that minimizes $\sum_{e\in p}(c_{e}-\mu_{e})$ and then subtracting the demand dual $\pi_{k}$. Since all edge weights $c_{e}\geq 0$ and $\mu_{e}\leq 0$, the modified costs $c_{e}-\mu_{e}\geq 0$, so the pricing problem reduces to a shortest path problem in a weighted graph.
@@ A.2 Runtime details
Instance Tree Path Source Objective
@@
grid1 0.0 0.0 0.0 0.0 0.0 0.6 8.2732e+05
grid2 0.0 0.1 0.0 0.0 0.0 0.6 1.7054e+06
grid3 0.0 0.0 0.0 0.0 0.3 0.8 1.5246e+06
grid4 0.0 0.1 0.0 0.0 0.5 0.8 3.0317e+06
grid5 0.0 0.1 0.1 0.1 1.8 1.1 5.0497e+06
grid6 0.3 0.1 0.2 0.1 2.6 1.4 1.0402e+07
grid7 0.4 0.1 0.3 0.1 12.1 2.8 2.5864e+07
grid8 1.6 0.2 1.3 0.2 101.3 2.9 4.1711e+07
grid9 3.4 0.2 2.8 0.2 46.8 4.0 8.2653e+07
grid10 4.3 0.3 4.1 0.3 53.3 4.4 1.6411e+08
grid11 3.9 0.2 4.4 0.3 58.1 4.4 3.2926e+08
grid12 4.4 0.2 4.2 0.3 175.1 8.1 5.7719e+08
grid13 10.2 0.3 12.4 0.4 200.2 8.8 1.1593e+09
grid14 5.5 0.3 8.2 0.4 457.8 16.7 1.8027e+09
grid15 10.6 0.4 18.8 0.6 531.2 15.4 3.5935e+09
@@
planar30 0.0 0.0 0.0 0.0 0.1 0.6 4.4351e+07
planar80 0.4 0.1 0.3 0.1 0.6 0.8 1.8244e+08
planar100 0.3 0.1 0.5 0.1 1.0 1.0 2.3134e+08
planar150 4.4 0.2 4.1 0.3 4.9 1.4 5.4809e+08
planar300 1.6 0.2 1.9 0.2 26.0 3.3 6.8998e+08
planar500 0.7 0.2 1.0 0.2 65.3 4.5 4.8198e+08
planar800 5.5 0.3 8.6 0.4 228.0 8.5 1.1674e+09
planar1000 75.3 0.6 53.5 0.9 426.2 25.7 3.4496e+09
planar2500 5894.7 10.0 866.9 19.3 - m 1.2662e+10
@@
Austin 670.6 14.5 4967.0 48.1 - m 5.4329e+06
Barcelona 0.1 0.1 2.5 0.2 24.8 2.7 8.8588e+01
BerlinCenter 10.3 0.5 19.0 0.7 - m 2.7529e+07
Birmingham 196.2 3.3 399.3 16.7 - m 2.1565e+05
ChicagoRegional 68.2 6.1 981.8 33.4 - m 5.5131e+06
ChicagoSketch 0.5 0.2 3.1 0.6 34.4 4.1 6.7059e+06
Philadelphia 380.6 7.4 1471.2 17.3 - m 2.5578e+07
Sydney 66.4 11.1 2520.9 59.7 - m 4.9653e+06
Winnipeg 0.0 0.1 2.9 0.2 18.9 3.1 2.3767e+02